Grade 12th passAlgebra

How to solve following question. a/x-1 + b/x-2 + c/x-3 + d/x-4 = 1 cannot have imaginary roots if a,b,c,d are real number of same sign.

Suyash Karn
9 Years agoGrade 12th pass

1 Answer

krunal
9 Years ago
(a/x-1)+(b/x-1)+(c/x-1)=1
on solving we get,
a+B+c+d-(X+2X+3x+4x)=x
a+B+c+d=11x
ax+bx+cx+dx=11x2.
o solving u can get quadratic with coeff of a=11 c=o and b=a+b+c+d.